OpenStudy (anonymous):

A sweater costs $15 and a scarf costs $5. You pay $105 for 9 sweaters and scarves combined. How many of each item did you purchase? Help!!! Fan and Medal

OpenStudy (anonymous):

You need to set it up as a equation first

OpenStudy (acxbox22):

2 equations as a matter of fact

OpenStudy (anonymous):

so i already know that s could represent sweaters and f could represent scarves

OpenStudy (anonymous):

now how would i set the equations up.

OpenStudy (anonymous):

You have a total of what u paid in all which is 105

OpenStudy (anonymous):

so 15s+5f=105?

OpenStudy (anonymous):

You cant use that equation cause they 9 also comes into play some how, try bumping the question lol sorry I thought i knew how but I dont.

OpenStudy (anonymous):

The guy earlier said you need to equtns. and he's right

OpenStudy (anonymous):

i know.

OpenStudy (anonymous):

sorry lol

OpenStudy (anonymous):

@acxbox22

OpenStudy (anonymous):

@Purpleloverr its ok

OpenStudy (acxbox22):

let x=sweaters and y=scarves x+y=9 since both combined she bought 9 and 15x+5y=105 to represent the costs and total cost

OpenStudy (anonymous):

ok. I know how to do the rest. Its just the equations that made me confused thanks.

OpenStudy (acxbox22):

np :)